Bicycling, the world's leading cycling magazine, reviews hundreds of bicycles throughout the year. Their "Road-Race" category contains reviews of bikes used by riders primarily interested in racing. One of the most important factors in selecting a bike for racing is the weight of the bike. The following data show the weight (pounds) and price ($) for 10 racing bikes reviewed by the magazine.†

Brand Weight Price ($)
FELT F5 17.8 2,100
PINARELLO Paris 16.1 6,250
ORBEA Orca GDR 14.9 8,370
EDDY MERCKX EMX-7 15.9 6,200
BH RC1 Ultegra 17.2 4,000
BH Ultralight 386 13.1 8,600
CERVELO S5 Team 16.2 6,000
GIANT TCR Advanced 2 17.1 2,580
WILIER TRIESTINA Gran Turismo 17.6 3,400
SPECIALIZED S-works Amira SL4 14.1 8,000

a) Predict the price (in dollars) for a bike that weighs 17 pounds. (Round your answer to the nearest dollar.)

b) Predict the price (in dollars) for a bike that weighs 15 pounds. (Round your answer to the nearest dollar.)
Answer #1

Solution:

a) Price for a bike that weights 17 pounds = 28574 - 1439(17)

= 28574 - 24463

= 4111

b) Price for a bike that weights 15 pounds = 28574 - 1439(15)

= 28574 - 21585

= 6989

Please up vote the solution if it helpful...........Thank you